# Multi-Rank Subspace Change-Point Detection for Monitoring Robotic Swarms

### v1

@@ 3.2 Multi-rank Subspace-CUSUM procedure
In practice, the exact CUSUM requires a priori knowledge of the post-change subspace $U$ and signal strengths ${\lambda_{i}}$, which is often unavailable. A straightforward generalized likelihood ratio approach would involve continuously re-estimating these parameters and plugging them into the likelihood ratio, but doing so naively at every time $t$ would be computationally intensive. Instead, we propose the multi-rank Subspace-CUSUM procedure, which tracks an estimated signal subspace on the fly and uses it to form a scalable detection statistic.
@@
We maintain an estimate $\hat{U}_{t}\in\mathbb{R}^{k\times d}$ of the leading $d$-dimensional signal subspace at time $t$. This can be obtained by computing the sample covariance over a recent sliding window of length $w$ and extracting its $d$ leading eigenvectors. In other words, at each time $t$ we form
$\hat{\Sigma}_{t}=\frac{1}{w}\sum_{i=t-w+1}^{t}x_{i}x_{i}^{T},$
and let $\hat{U}_{t}=[\hat{u}_{1,t},\dots,\hat{u}_{d,t}]\in\mathbb{R}^{k\times d}$ be the matrix of $d$ principal unit norm eigenvectors $\hat{u}_{i,t}$, $i=1,\dots,d$, of $\hat{\Sigma}_{t}$, corresponding to its $d$ largest eigenvalues. This $\hat{U}_{t}$ serves as an approximation to the true $U$ under the post-change regime. The window size $w$ should be large enough to reliably learn the subspace but not so large that it fails to adapt when the covariance shifts. In the initial pre-change period, $\hat{U}_{t}$ will essentially track arbitrary directions in noise, but once the covariance spikes appear, $\hat{U}_{t}$ will begin to align with the true signal subspace $U$ after on the order of $w$ samples. Thus, $w$ can be viewed as an estimation lag that contributes to the overall detection delay.
@@
Given a delayed subspace estimate $\hat{U}_{t+w}$ computed over time $[t+1,t+w]$ (practically, this is always possible by properly delaying our data by $w$ samples), we define a scalar increment $Z_{t}$ that measures how signal-like the new observation $x_{t}$ is. A natural choice is the squared norm of $x_{t}$ projected onto the estimated subspace:
$Z_{t}=\lVert\hat{U}_{t+w}^{T}x_{t}\rVert^{2}=\sum_{i=1}^{d}(\hat{u}_{i,t+w}^{T}x_{t})^{2}.$
@@
Here, $Z_{t}$ represents the energy of $x_{t}$ in the top-$d$ principal directions estimated from future $w$ data.
We define the Subspace-CUSUM statistic $S_{t}^{\text{sub}}$ recursively as
$S_{t}^{\text{sub}}=(S_{t-1}^{\text{sub}})^{+}+Z_{t}-\Delta,\quad S_{0}^{\text{sub}}=0.$
@@
The drift parameter $\Delta$ should be set between the typical values of $Z_{t}$ under the no-change and change scenarios. Formally, we require the double inequality
$\mathbb{E}_{\infty}[Z_{t}]<\Delta<\mathbb{E}_{0}[Z_{t}],$
@@
so that $S_{t}^{\text{sub}}$ has a negative drift before the change and a positive drift after the change. The stopping time is
$T^{\text{sub}}=\inf\{t:S_{t}^{\text{sub}}\geq b\}$
for some threshold $b$ calibrated to achieve the desired false alarm rate.
@@
The multi-rank Subspace-CUSUM procedure involves a trade-off between estimation accuracy and detection speed. Before the procedure can reliably raise an alarm, the subspace estimate $\hat{U}_{t}$ must capture the new spike directions. This estimation window of length $w$ contributes $w$ samples of delay after the true change-point, during which $\hat{U}_{t}$ is converging to $U$ and $S_{t}^{\text{sub}}$ may not yet be growing. Once $\hat{U}_{t}$ successfully approximates $U$, the statistic $S_{t}^{\text{sub}}$ sees increments $Z_{t}-\Delta$ with a positive mean, and thereafter $S_{t}^{\text{sub}}$ behaves similarly to the exact CUSUM.
4 Analysis
@@ 4.1 Drift parameter
In this section, we analyse the drift parameter $\Delta$ of the detection statistic to the rank-$d$ spiked covariance model. Recall that in the Subspace-CUSUM procedure, at each time $t$ the statistic is updated by adding the quantity $Z_{t}-\Delta$, with a reset to zero if the sum becomes negative. The main challenge is to characterize $\mathbb{E}_{\infty}Z_{t}$ and $\mathbb{E}_{0}Z_{t}$ in order to pick an appropriate $\Delta$. We address this by analyzing the increment term $Z_{t}$ under each regime.
Importantly, by our procedure design, $\hat{U}_{t+w}$ is computed from a moving window that does not include the current sample $x_{t}$. This assures the independence of $\hat{U}_{t+w}$ and $x_{t}$, which greatly simplifies the analysis. That is,
@@
Thus, under pre-change distribution, the expected increment is $d\sigma^{2}$.
@@
To specify the post-change increment, we invoke the Central Limit Theorem argument for the eigenvectors of the sample covariance matrix $\hat{\Sigma}_{t}$. Assuming that the population eigenvalues $\{\lambda_{i}:i=1,\dots,d\}$ are distinct, each sample eigenvector $\hat{u}_{i}$ is approximately Gaussian distributed around the true eigenvector $u_{i}$, with covariance scaling as $w^{-1}$ (paul2007asymptotics, ). Based on this approximation, we derive the following expression for the post-change mean of the increment. A detailed derivation can be found in Appendix A.
@@
Assume that $\{\lambda_{i}:i=1,\dots,d\}$ are distinct. Under the post-change regime, we have the following mean values of increments: As $w\to\infty$,
$\mathbb{E}_{0}[Z_{t}]=\sigma^{2}\sum_{i=1}^{d}(1+\rho_{i})+\frac{\sigma^{2}}{w}\sum_{i=1}^{d}(1+\rho_{i})\Big{[}\frac{k-1}{\rho_{i}^{2}}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}\frac{(1+\rho_{j})^{2}}{(\rho_{i}-\rho_{j})^{2}}\Big{]}+o\big{(}\frac{1}{w}\big{)}.$ (6)
In other words, with the choice of large $w$, the post-change increment is about $d$ times the noise variance plus the total signal power in the $d$ spike components.
Combining the above results, we find that in the rank-$d$ spiked covariance model,
  • Pre-change increment: $\mathbb{E}_{\infty}Z_{t}=d\sigma^{2}$.
  • Post-change increment: $\mathbb{E}_{0}Z_{t}=d\sigma^{2}+\sum_{i=1}^{d}\lambda_{i}+O(1/w)$.
As long as at least one $\lambda_{i}>0$, we have $\lim_{w\to\infty}\mathbb{E}_{0}Z_{t}>\mathbb{E}_{\infty}Z_{t}$. Thus, there is a strictly positive gap between the post-change and pre-change means of the statistic’s increment, which allows us to choose a drift $\Delta$ in between these values.
In practice, one may not know the exact ${\lambda_{i}}$ in advance. However, we can design $\Delta$ based on a conservative minimum signal strength. For example, suppose it is known that each signal component has SNR at least $\rho_{\min}=\lambda_{\min}/\sigma^{2}$. Then a safe choice is to set $\Delta$ to the midpoint between the no-signal and minimal-signal expectations:
@@
This choice guarantees $\mathbb{E}_{\infty}Z_{t}=d\sigma^{2}<\Delta<d\sigma^{2}+d\lambda_{\min}\leq\lim_{w\to\infty}\mathbb{E}_{0}Z_{t}$, so if $w$ is sufficiently larger than $k$, the required inequality is satisfied even in the worst-case scenario when each signal component is uniformly at the minimum strength.

### v2

@@ 3.2 The MRS-C procedure
In practice, the Exact CUSUM requires a priori knowledge of the post-change subspace $U$ and signal strengths $\lambda_{i}$}, which is often unavailable. A straightforward generalized likelihood ratio approach would involve continuously re-estimating these parameters and plugging them into the likelihood ratio, but doing so naively at every time $t$ would be computationally intensive. Instead, we propose the Multi-rank Subspace-CUSUM (MRS-C) procedure, which tracks an estimated signal subspace in real time and uses it to construct a scalable detection statistic.
@@
We maintain an estimate $\hat{U}_{t}\in\mathbb{R}^{k\times d}$ of the leading $d$-dimensional signal subspace at time $t$. At each time step, this estimate is obtained by forming the sample covariance matrix over a sliding window of length $w$,
$\hat{\Sigma}_{t}=\frac{1}{w}\sum_{i=t-w+1}^{t}x_{i}x_{i}^{\top},$
and extracting its $d$ leading eigenvectors. Specifically, we let $\hat{U}_{t}=[\hat{u}_{1,t},\dots,\hat{u}_{d,t}]$ denote the matrix of unit-norm eigenvectors of $\hat{\Sigma}_{t}$ corresponding to its $d$ largest eigenvalues. Under the post-change regime, $\hat{U}_{t}$ provides a plug-in estimate of the true signal subspace $U$.
The window length $w$ controls the trade-off between estimation accuracy and adaptivity. Larger values of $w$ improve subspace estimation but delay responsiveness to changes in the covariance structure. Before the change, the estimate reflects only noise fluctuations, whereas after the appearance of a low-rank spike, it progressively aligns with $U$. Accordingly, the choice of $w$ directly affects the detection delay and must be balanced against the desired false-alarm performance.
When estimating the signal subspace at each time step, we deliberately avoid using past observations and instead construct the estimate from the next $w$ samples. This future-window strategy has a key advantage that it ensures that the subspace estimate formed from $\{x_{t+1},\ldots,x_{t+w}\}$ is independent of the current observation $x_{t}$, which greatly simplifies the drift calculations and theoretical analysis presented later.
@@
Given such a delayed subspace estimate $\hat{U}_{t+w}$ computed over the time interval $[t+1,t+w]$, we quantify how “signal-like” the current observation $x_{t}$ is through a scalar increment $Z_{t}$. A natural choice is the squared norm of the projection of $x_{t}$ onto the estimated subspace,
$Z_{t}=\big\|\hat{U}_{t+w}^{T}x_{t}\big\|^{2}=\sum_{i=1}^{d}\big(\hat{u}_{i,t+w}^{T}x_{t}\big)^{2}.$ (6)
@@
Here, $Z_{t}$ represents the energy of $x_{t}$ in the top-$d$ principal directions inferred from the future window of $w$ observations.
Accordingly, we define the MRS-C statistic $S_{t}^{\mathrm{sub}}$ recursively as
$S_{t}^{\mathrm{sub}}=(S_{t-1}^{\mathrm{sub}})^{+}+Z_{t}-\Delta,\quad S_{0}^{\mathrm{sub}}=0.$ (7)
@@
The drift parameter $\Delta$ should be set between the typical values of $Z_{t}$ under the pre-change and post-change scenarios. Formally, we require the double inequality
$\mathbb{E}_{\infty}[Z_{t}]<\Delta<\mathbb{E}_{0}[Z_{t}],$ (8)
@@
so that $S_{t}^{\mathrm{sub}}$ has a negative drift before the change and a positive drift after the change. Selection of the drift will be discussed rigorously in Section 4.1. The detection time is
$T^{\mathrm{sub}}=\inf\{t:S_{t}^{\mathrm{sub}}\geq b\}+w$ (9)
for a threshold $b$ calibrated via Monte-Carlo simulation to achieve the desired false alarm rate. The extra “+w+w” term reflects the fact that the statistic at time $t$ is based on a subspace estimated from the future window of length $w$.
@@
The MRS-C procedure, therefore, involves a trade-off between subspace-estimation accuracy and detection speed. Before the procedure can reliably raise an alarm, the delayed estimate $\hat{U}_{t}$ must capture the new spike directions. This estimation window introduces an inherent delay after the true change-point during which $\hat{U}_{t}$ is converging to $U$ and $S_{t}^{\mathrm{sub}}$ may not yet be growing. Once $\hat{U}_{t}$ sufficiently approximates $U$, the statistic adds increments $Z_{t}-\Delta$ with positive mean, and thereafter $S_{t}^{\mathrm{sub}}$ behaves similarly to the Exact CUSUM.
4 Theoretical analysis
This section develops the theoretical performance guarantees for the proposed MRS-C procedure under the emerging subspace model. We first characterize the mean behavior of the detection statistic before and after the change, which leads to principled choices of the drift parameter and the window length. Building on these results, we derive asymptotic expressions for the EDD under a prescribed ARL constraint and establish first-order asymptotic optimality relative to the oracle Exact CUSUM.
@@ 4.1 Drift parameter
In this subsection, we analyse the drift parameter $\Delta$ of the detection statistic under the rank-$d$ spiked covariance model. Recall that the statistic is updated at each time step by adding the quantity $Z_{t}-\Delta$, with a reset to zero whenever the cumulative sum becomes negative. To ensure correct operation of the procedure, $\Delta$ must lie strictly between the pre-change and post-change expectations of $Z_{t}$. Accordingly, our main task is to characterize $E_{\infty}[Z_{t}]$ and $E_{0}[Z_{t}]$ under the two regimes.
Importantly, by our procedure design, the subspace estimate $\hat{U}_{t+w}$ is computed from a moving window that does not include the current sample $x_{t}$. This ensures the independence of $\hat{U}_{t+w}$ and $x_{t}$, which allows us to simplify the expectation as
@@
To specify the post-change increment, we utilize the asymptotic properties of the sample eigenvectors. Throughout Section 4, we impose the following assumption for theoretical analysis.
Assumption 4.1.
Assume the following for the true subspace covariance structure.
  1. (i)
    The subspace dimension $d$ remains a constant.
  2. (ii)
    The population eigenvalues $\{\lambda_{i}:i=1,\dots,d\}$ are distinct and are ordered as $\lambda_{1}>\dots>\lambda_{d}>0$.
Under Assumption 4.1, each sample eigenvector $\hat{u}_{i,t+w}$ is approximately Gaussian distributed around the true eigenvector $u_{i}$, with covariance scaling as $w^{-1}$ (anderson1963asymptotic; paul2007asymptotics). Based on this approximation, we derive the following expression for the post-change mean of the increment. A detailed derivation can be found in Appendix A.
Remark 4.2 (Eigenvalue multiplicity).
Assumption 4.1(ii) requires distinct spike eigenvalues. This condition is imposed for proof convenience in the eigenvector perturbation arguments. The proposed detection statistic depends only on the estimated rank-$d$ signal subspace through the projection matrix, rather than on the ordering of individual eigenvectors. Consequently, the procedure remains well-defined when the spikes have multiplicity, as in several of our simulation settings in Section 6.
@@
Under Assumption 4.1, the expected increment under the post-change regime is given by, as $w\to\infty$,
$\mathbb{E}_{0}[Z_{t}]=\sigma^{2}\sum_{i=1}^{d}(1+\rho_{i})-\frac{(k-d)\sigma^{2}}{w}\sum_{i=1}^{d}\frac{1+\rho_{i}}{\rho_{i}}+o\Big(\frac{1}{w}\Big).$ (11)
Again, for the MRS-C procedure to function as intended, the drift parameter $\Delta$ must satisfy the condition (8). Comparing the expectations, we can specify a condition for the window size $w$ to ensure $\mathbb{E}_{0}[Z_{t}]>\mathbb{E}_{\infty}[Z_{t}]$ by requiring the signal term to dominate the estimation error term:
$w>\frac{k-d}{\sum_{i=1}^{d}\rho_{i}}\sum_{i=1}^{d}\frac{1+\rho_{i}}{\rho_{i}}.$
Under this condition, $\Delta$ can take any value in the valid range.
In practice, the exact signal strengths $\{\lambda_{i}\}$ may not be known in advance. However, we can design $\Delta$ based on a conservative lower bound for the signal strength. For example, suppose it is known that each signal component has SNR at least $\rho_{\min}=\lambda_{\min}/\sigma^{2}$. A robust choice is to set $\Delta$ to the midpoint between the pre-change expectation and a conservative estimate of the post-change expectation as
@@
This choice guarantees $\mathbb{E}_{\infty}Z_{t}<\Delta$ immediately. Furthermore, provided that $w$ is chosen sufficiently large to satisfy the condition above, it also ensures $\Delta<\mathbb{E}_{0}Z_{t}$, thereby enabling effective detection even in the worst-case scenario where signal strengths are minimal.
@@
4.2 Asymptotic EDD and oracle efficiency
In this section, we establish the asymptotic optimality properties of the proposed MRS-C procedure. Building on the drift analysis in Section 4.1, we derive asymptotic expressions for the EDD as the ARL tends to infinity. We then identify asymptotically optimal choices of the threshold $b$, drift $\Delta$, and window length $w$ that minimize the detection delay, and show that the resulting performance matches that of the oracle Exact CUSUM up to a multiplicative efficiency constant.
The following Proposition 4.5 provides an explicit asymptotic characterization of the EDD for the MRS-C procedure. The result follows the classical nonlinear renewal framework for sequential change-point procedures (tartakovsky2014sequential, Chapter 3).
Proposition 4.5 (First-order optimal EDD for arbitrary window length).
For a prescribed ARL level $\mathbb{E}_{\infty}[T^{\mathrm{sub}}]=\gamma>1$, as $\gamma\to\infty$, if we select the threshold $b$ and the drift $\Delta$ as
$b$ $=\frac{2\sigma^{2}\log\gamma(1+o(1))}{1-\frac{d}{A}},$ (12)
$\Delta$ $=\frac{d\sigma^{2}}{1-\frac{d}{A}}\log(\frac{A}{d}),$ (13)
where $A=A(w)$ represents the dominant term of the post-change increment expectation as
$A\coloneqq\sum_{i=1}^{d}(1+\rho_{i})\Big(1-\frac{k-d}{w\rho_{i}}\Big),$ (14)
then the EDD under the alternative is given by
$\mathbb{E}_{0}[T^{\mathrm{sub}}]=\frac{2\log\gamma(1+o(1))}{A-d(1+\log(\frac{A}{d}))}+w.$ (15)
In particular, Proposition 4.5 identifies a pair $(b,\Delta)$ that asymptotically equalizes the exponential overshoot distribution under the pre-change measure. The condition $\mathbb{E}_{\infty}\!\left[e^{\delta_{\infty}(Z_{t}-\Delta)}\right]=1$ determines the so-called adjustment coefficient $\delta_{\infty}$, a standard object in nonlinear renewal theory, ensuring that the stopping rule behaves like a first-order optimal CUSUM under the nominal distribution. The detailed proof is given in Appendix A.
The expression for $A$ in (14) represents the leading-order post-change drift of the statistic, incorporating both the spike strengths $\{\rho_{i}\}$ and the window-induced eigenvector estimation error at order $1/w$. The asymptotically optimal delay in (15) is decomposed into two components: the intrinsic information-theoretic delay, and the estimation lag $w$ due to the sliding window needed to estimate the evolving subspace independently. The intrinsic delay scales logarithmically with the ARL, as in the classical CUSUM theory.
Minimizing the EDD expression in Proposition 4.5 with respect to the window length $w$ yields the optimal parameters.
Corollary 4.6 (Optimal parameters).
To minimize the expected detection delay, the asymptotically optimal window size $w^{*}$ is given by
$w^{*}=\sqrt{\log\gamma}\cdot\frac{\sqrt{2(k-d)(\sum_{i=1}^{d}\frac{1+\rho_{i}}{\rho_{i}})\Big(\frac{\sum_{i=1}^{d}\rho_{i}}{\sum_{i=1}^{d}(1+\rho_{i})}\Big)}}{\sum_{i=1}^{d}\rho_{i}-d\log(\frac{1}{d}\sum_{i=1}^{d}(1+\rho_{i}))}(1+o(1)).$ (16)
Hence, corresponding optimal choice of the threshold $b$ and the drift $\Delta$ are given by
$b\asymp\frac{2\sigma^{2}\log\gamma}{1-\frac{d}{A(w^{*})}},\qquad\Delta=\frac{d\sigma^{2}}{1-\frac{d}{A(w^{*})}}\log(\frac{A(w^{*})}{d}).$ (17)
Finally, we establish the efficiency of our procedure relative to the oracle Exact CUSUM.
Theorem 4.7 (First-order asymptotic optimality).
The ratio of the expected detection delays of the MRS-C ($T^{\mathrm{sub}}$) to the oracle Exact CUSUM ($T_{C}$) is given by
$\frac{\mathbb{E}_{0}[T^{\mathrm{sub}}]}{\mathbb{E}_{0}[T_{C}]}=\mathcal{K}+O\Big(\frac{1}{\sqrt{\log\gamma}}\Big),$ (18)
where the efficiency constant $\mathcal{K}\geq 1$ is defined as
$\mathcal{K}=\frac{\sum_{i=1}^{d}[\rho_{i}-\log(1+\rho_{i})]}{\sum_{i=1}^{d}[\rho_{i}-\log(1+\bar{\rho})]}\geq 1,$ (19)
with $\bar{\rho}=\frac{1}{d}\sum_{i=1}^{d}\rho_{i}$ representing the average signal-to-noise ratio. The constant $\mathcal{K}$ attains its minimum value of 1 if and only if all signal strengths are uniform (i.e., $\rho_{i}=\bar{\rho}$ for all $i$).
Remark 4.8 (Difference from rank-one case).
Our theoretical results generalize and unify the findings from prior studies on subspace and network change detection. First, xie2020sequential; xie2018first address the rank-1 spiked covariance model. Our multi-rank framework naturally encompasses their result as a special case. When $d=1$, the efficiency constant $\mathcal{K}$ trivially becomes 1, recovering their finding that subspace tracking achieves full oracle efficiency asymptotically in the rank-1 setting. However, for $d>1$, our Theorem 4.7 reveals that heterogeneity in signal strengths $\{\rho_{i}\}$ introduces a penalty factor $\mathcal{K}>1$. Compared to zhang2023spectral, which deals with inverse covariance structure with adjacency matrices, despite the different statistical models, both methods exhibit a similar optimal window scaling of $w^{*}\asymp\sqrt{\log\gamma}$, suggesting a universal principle for window-limited CUSUM procedures involving subspace estimation. However, the specific drift terms and the efficiency constant $\mathcal{K}$ are unique to the geometry of the spiked covariance model.
5 Parallel procedure for dimension selection
Throughout our discussion and analyses thus far, we have assumed that the dimension $d$ of the underlying subspace is known. In practice, however, the true subspace dimension is typically unknown and must be estimated from data or chosen adaptively. To handle the unknown $d$, we now consider running multiple detection procedures in parallel across a set of candidate dimensions $\mathcal{D}=\{d_{1},\dots,d_{m}\}$. Motivated by xie2013sequential, by setting detection thresholds for each dimension individually and ensuring the overall false-alarm rate is properly controlled, the earliest detection from these parallel runs can be taken as the final decision.
Formally, for a given candidate dimension $d_{k}\in\mathcal{D}$, $k=1,\dots,m$, the single-procedure update is
$Z_{t}^{(d_{k})}=\sum_{i=1}^{d_{k}}\big(\hat{u}_{i,t+w}^{T}x_{t}\big)^{2},\quad S_{t}^{(d_{k})}=(S_{t-1}^{(d_{k})})^{+}+Z_{t}^{(d_{k})}-d_{k}\Delta_{1},$
with $S_{0}^{(d_{k})}=0$, and in our implementation we set $\Delta_{1}=\sigma^{2}\big(1+\rho_{\min}/2\big)$, following the guideline from Section 4. Each $S_{t}^{(d_{k})}$ is compared against a threshold $b^{(d_{k})}$, where each threshold $b^{(d_{k})}>0$ is chosen to achieve the desired ARL under the pre-change distribution. We declare that the procedure associated with dimension $d_{k}$ signals a change at time
$T^{(d_{k})}=\inf\{t:S_{t}^{(d_{k})}\geq b^{(d_{k})}\}.$
The parallel stopping rule for the overall procedure is the minimum of these stopping times all candidate dimensions
$T_{\text{mix}}=\min_{d_{k}\in\mathcal{D}}T^{(d_{k})}=\min_{k\in[m]}\inf\{t:S_{t}^{(d_{k})}\geq b^{(d_{k})}\}.$ (20)
In words, $T_{\text{mix}}$ is the first time any one of the $m$ parallel statistics signals a change. The index achieving the minimum in (20), $k^{*}=\mathop{\rm argmin}_{k}T^{(d_{k})}$, effectively suggests an estimate of the post-change subspace dimension as
$\hat{d}=d_{k^{*}}.$ (21)
The rationale is that whichever candidate dimension $d_{k}$ best matches the true subspace will accumulate evidence the fastest.
Under the fixed threshold regime, given the target ARL $\gamma$, we employ a simple and conservative Bonferroni-type union bound using the fact that
$\mathbb{P}_{\infty}(T_{\mathrm{mix}}\leq t)\leq\sum_{k=1}^{m}\mathbb{P}_{\infty}(T^{(d_{k})}\leq t).$
Therefore, each threshold $b^{(d_{k})}$ shall be calibrated to have ARL $m\cdot\gamma$ by Monte-Carlo simulation.